While on a walk in the country, you pass a field full of goats and chickens. After a quick count, you determine there are
31 heads and 86 feet in the field. How many of each animal are there?

Found 3 solutions by Boreal, MathTherapy, ikleyn:
Answer by Boreal(15235)

G+C=31, where G=goats and C=chickens. This is the count equation.
4G+2C=86; this is the feet equation.
-2G-2C=-62
2G=24
G=12---48 feet
C=10--38 feet
They add to 86 feet
(8x+4=52)
x=6

Answer by MathTherapy(10555)


While on a walk in the country, you pass a field full of goats and chickens. After a quick count, you determine there are
31 heads and 86 feet in the field. How many of each animal are there?
 


Answer by ikleyn(52832)

.
See the lesson
    - Problem on animals at a farm
in this site.

Also, you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lesson is the part of this online textbook under the topic "Miscellaneous word problems".